### v1

@@ 3.3 BLOCKCHAIN-BASED IDENTIFICATION AND AUTHORIZATION PROTOCOL
The detailed algorithm is provided in Algorithm 1, which formalizes the identification and authorization steps between communicating nodes.
Figure 6: Zero-Trust Decentralized Blockchain-based Communication Example Scenario for Autonomous Vehicles
Performance considerations include the added overhead of blockchain communication and consensus. Lightweight permissioned blockchains and optimized consensus protocols (e.g., Hyperledger Iroha) are used to reduce latency, making the system suitable for urban and highway deployment scenarios. Detailed performance metrics are presented in Section 4.
@@
In summary, the proposed D-IM system leverages blockchain and ZTA principles to enable decentralized, tamper-proof, and continuously verifiable identity management in V2X environments. By ensuring rigorous authentication, efficient key establishment, and protection against classical vehicular attacks, the framework provides a strong security foundation for autonomous V2X networks while remaining scalable for real-world deployment. Future work will investigate the integration of smart contracts and anomaly-based ML detection to automate misbehavior penalties and enhance the adaptability of framework.
@@
$p_{k}$ (resource retention) 0 or 0.8
Sensing threshold -110 dBm
Figure 7: (a) Highway Packet Reception Rate and (b) Urban Packet Reception Rate vs. the distance using LTE-V2X RAT and average packet size required by our implementation.
4 EXPERIMENTAL EVALUATION
Figure 6 presents a representative scenario of the framework architecture in action. Imagine a busy intersection where vehicles and roadside units (RSUs) are constantly exchanging information to maintain smooth traffic flow. Each vehicle and RSU carries its own copy of the framework code, acting as a self-contained guardian that enforces the principles of zero trust. As cars approach the intersection, they broadcast their intentions and receive updates from others. At the same time, RSUs coordinate signal changes, all while communicating through the blockchain network.
In this scenario, the blockchain serves as the trusted backbone, ensuring that communication remains secure and that vehicle identities are managed in a decentralized and tamper-resistant manner. But the system goes further: embedded within each node is an anomaly behavior analysis engine that continuously evaluates the plausibility of exchanged messages. If, for example, one car suddenly reports a location that conflicts with its prior GPS path, or if a traffic light sends inconsistent signals, the framework recognizes these as potential GPS spoofing or fake data injection attempts. Such anomalies are not handled in isolation, nodes collaborate to resolve conflicts through decentralized consensus, preventing malicious actors from undermining the system.
Policies and access controls further refine this process, dictating which entities are authorized to transmit or act on certain types of data. In this way, the architecture creates a layered shield, where local enforcement, blockchain-backed consensus (Left for future work), and multi-level anomaly detection (See our work in [27] together preserve the integrity of vehicular communication. This scenario illustrates not only how the system reacts to ordinary coordination needs but also how it adapts and defends itself when adversarial conditions arise, thereby highlighting the resilience and practicality of the framework. The following subsections dive into the D-IM system implementation details and the simulations carried out to evaluate it.
4.1 BLOCKCHAIN-BASED IDENTIFICATION PROTOCOL IMPLEMENTATION
This subsection describes the implementation of the blockchain network and the proposed identification protocol. For the blockchain infrastructure, we adopted Hyperledger Iroha v2, a permissioned blockchain platform designed for ease of integration and identity-centric use cases [86]. Iroha emphasizes simplicity, modularity, and reliability, making it well-suited for distributed ledger technology (DLT) applications such as identity management and asset tracking. Its lightweight architecture and support for multiple programming languages provide flexibility for rapid prototyping and deployment.
@@ 4.1 BLOCKCHAIN-BASED IDENTIFICATION PROTOCOL IMPLEMENTATION
Consensus within Iroha is achieved through the Yet Another Consensus (YAC) protocol [87], a crash-fault-tolerant variant of Byzantine Fault Tolerance (BFT). By focusing on high throughput and low latency rather than full Byzantine resilience, YAC achieves eventual consistency across network nodes while maintaining system resilience in the event of node failures. This efficiency is especially beneficial for time-sensitive vehicular scenarios where transaction confirmation speed is critical. This makes Iroha an efficient choice for consortium blockchain deployments where participants are semi-trusted, striking a balance between resilience and performance. While Hyperledger Fabric is often adopted for complex blockchain applications, Iroha’s lightweight design and rapid development cycle make it a strong candidate for vehicular identity management.
@@
The proposed identification and authorization protocol was implemented in Python. Communication with the Iroha blockchain was facilitated through the official Iroha Python library v0.0.5.5. Cryptographic operations were realized using the cryptography package v3.4.8, which provides RSA and symmetric key primitives, while Python’s secrets module generated random secret values for Diffie–Hellman key exchange. For vehicle-to-vehicle (V2V) communication, we employed the Python socket module to establish and manage encrypted communication channels between participants.
Figure 8: (a) Highway Packet Reception Rate and (b) Urban Packet Reception Rate vs. Number of Vehicles/km at 150m.
Figure 9: (a) Highway Channel Busy ratio and (b) Urban Packet Reception Rate vs. number of vehicles/km
4.2 SIMULATION TOOLS AND SETTINGS
@@
To evaluate the proposed protocol, we used WiLabV2XSim v6.2 a discrete-event simulation framework developed in MATLAB for modeling vehicular networks with a focus on cooperative awareness services. WiLabV2XSim supports both DSRC and C-V2X communications and can incorporate synthetic mobility models or real traffic traces to represent vehicle movements [21].
Figure 10: (a) The Highway Data Age and (b) The Urban Data Age vs. number of vehicles/km
The main simulation parameters are summarized in Table 2. Vehicles broadcast messages at fixed intervals of 100 ms (10 Hz) with a beacon size of either 90 B (minimum packet size exchanged in our identification protocol implementation) or 1670 B (average packet size exchanged in our identification protocol implementation). All nodes use the same transmission power of 23 dBm and antennas with 3 dB gain. The propagation environment is modeled using the WINNER+ model with a path loss exponent of 4, and log-normal shadowing with standard deviations of 3 dB (LOS) and 4 dB (NLOS). For DSRC, channel access is modeled using carrier-sense multiple access with collision avoidance (CSMA/CA), while C-V2X employs sensing-based semi-persistent scheduling (SPS).
The framework was evaluated using 3 output parameters:

### v2

@@ 4 DECENTRALIZED IDENTITY MANAGEMENT (D-IM) SYSTEM ARCHITECTURE
    • Step 12: Secure data exchange begins using $K_{s}$
5 SECURITY ANALYSIS
5.1 Formal BAN Logic-Based Security Analysis
To formally demonstrate the security properties of the proposed Zero Trust-based Decentralized Identity Management, D-IM, system, we employ Burrows-Abadi-Needham (BAN) logic [UCAM-CL-TR-138, burrows1990logic] to provide a formal proof that our system goals have been achieved. BAN logic is a formal security hypothesis based security analysis used to prove the validity of securiry assumptions. BAN logic has been extensively applied to validate mutual authentication, key agreement, and freshness guarantees in security-critical systems, including vehicular and wireless architectures [abdel2022proxy, farrea2025zero, guo2024uava]. BAN logic allows us to model how two vehicles evolve their beliefs about each other’s identity and about the established session key through the exchange of protocol messages, under clearly specified assumptions.
Table 3 summarizes the BAN logic notations used throughout this analysis.
Table 3: BAN Logic Notations
Notation Description
$X\mid\equiv M$ Entity $X$ believes that statement $M$ is true
$X\triangleleft M$ Entity $X$ has received message $M$
$X\mid\sim M$ Entity $X$ once said message $M$
$X\Rightarrow M$ Entity $X$ has jurisdiction over statement $M$
#(M)\#(M) Message $M$ is fresh
$\leftrightarrow_{K}Y XX$ and $Y$ share a secret key $K$
$K_{pbX}$ Public key of entity $X$
$K_{pvX}$ Private key of entity $X$
The BAN logic inference rules applied in this work are listed in Table 4.
Table 4: BAN Logic Inference Rules
Rule Description
Message Meaning $\frac{X\mid\equiv K_{pbY},\quad X\triangleleft\{M\}_{K_{pvY}}}{X\mid\equiv Y\mid\sim M}$
Nonce Verification $\frac{X\mid\equiv\#(M),\quad X\mid\equiv Y\mid\sim M}{X\mid\equiv Y\mid\equiv M}$
Jurisdiction $\frac{X\mid\equiv Y\Rightarrow M,\quad X\mid\equiv Y\mid\equiv M}{X\mid\equiv M}$
In the D-IM setting, the main participants are Vehicle 1 ($V_{1}$) and Vehicle 2 ($V_{2}$), which mutually authenticate and establish a shared session key $K_{s}$ using the blockchain as a decentralized trust anchor. Blockchain verification peers (RSUs) maintain the ledger, but from the BAN perspective they are abstracted as a single trusted authority that certifies public keys and identity bindings. The analysis focuses on the identification and authorization protocol (Algorithm 1), where $V_{1}$ and $V_{2}$ verify each other’s credentials via the blockchain and then perform a Diffie-Hellman-based key exchange. We denote the blockchain authority (the set of verification peers / RSUs that maintain the ledger) as $BC$ for the purpose of initial trust assumptions.
5.1.1 Security Goals and Assumptions
In accordance with Zero Trust principles, the D-IM protocol does not rely on implicit trust and instead enforces continuous verification and explicit belief derivation. The following security goals are defined for the interaction between $V_{1}$ and $V_{2}$:
  • SG1: $V_{1}\mid\equiv V_{2}$
    Vehicle $V_{1}$ believes that it is communicating with a legitimate $V_{2}$.
  • SG2: $V_{2}\mid\equiv V_{1}$
    Vehicle $V_{2}$ believes that it is communicating with a legitimate $V_{1}$.
  • SG3: $V_{1}\mid\equiv V_{1}\leftrightarrow_{K_{s}}V_{2}$
    Vehicle $V_{1}$ believes that it shares a fresh session key $K_{s}$ with $V_{2}$.
  • SG4: $V_{2}\mid\equiv V_{1}\leftrightarrow_{K_{s}}V_{2}$
    Vehicle $V_{2}$ believes that it shares a fresh session key $K_{s}$ with $V_{1}$.
The following initial assumptions are derived from the design of the D-IM system:
  • IA1: $V_{1}\mid\equiv BC\Rightarrow(ID_{2},K_{pb2},Hash(ID_{2}||K_{pvt_{2}}))$
    $V_{1}$ believes that $BC$ has jurisdiction over $V_{2}$’s identity, public key, and corresponding hash.
  • IA2: $V_{2}\mid\equiv BC\Rightarrow(ID_{1},K_{pb1},Hash(ID_{1}||K_{pvt_{1}})))$
    $V_{2}$ believes that $BC$ has jurisdiction over $V_{1}$’s identity, public key, and corresponding hash.
  • IA3: $V_{1}\mid\equiv\#(Q)$
    $V_{1}$ believes that its Diffie-Hellman contribution $Q$ is fresh.
  • IA4: $V_{2}\mid\equiv\#(R)$
    $V_{2}$ believes that its Diffie-Hellman contribution $R$ is fresh.
  • IA5: $V_{1}\mid\equiv K_{pb2}$ is the authentic public key of $V_{2}$ since it got it from $BC$.
  • IA6: $V_{2}\mid\equiv K_{pb1}$ is the authentic public key of $V_{1}$ since it got it from $BC$.
  • IA7: $V_{1}\mid\equiv V_{2}\Rightarrow(V_{1}\leftrightarrow_{K_{s}}V_{2})$
    $V_{1}$ believes $V_{2}$ has jurisdiction over statements about the shared session key with $V_{1}$.
  • IA8: $V_{2}\mid\equiv V_{1}\Rightarrow(V_{1}\leftrightarrow_{K_{s}}V_{2})$
    $V_{2}$ believes $V_{1}$ has jurisdiction over statements about the shared session key with $V_{2}$.
These assumptions reflect that (i) public keys and identity bindings are certified by the blockchain, (ii) Diffie-Hellman contributions are fresh, and (iii) each vehicle trusts the other party to control statements about the shared session key.
5.1.2 D-IM BAN Logic-based Security Analysis
The core messages of our proposed identification and authorization protocol (Algorithm 1) are as follows:
  • Credential retrieval via blockchain
    $V_{1}$ $\triangleleft(ID_{2},K_{pb2},Hash(ID_{2}||K_{pvt_{2}})))\text{ from }BC,$
    $V_{2}$ $\triangleleft(ID_{1},K_{pb1},Hash(ID_{1}||K_{pvt_{1}})))\text{ from }BC.$
    After this step, using IA1–IA2 and the jurisdiction rule, both vehicles believe that the retrieved identity, public key and hash information are correct.
  • Vehicle 1 $\rightarrow$ Vehicle 2: $M_{1}:V_{2}\triangleleft\{ID_{2},K_{pb1},TS,Q,Hash(ID_{2}||K_{pvt_{2}}))\}_{K_{pb2}}$ $V_{1}$ sends its Diffie–Hellman value $Q$ and related information encrypted with $V_{2}$’s public key.
  • Vehicle 2 $\rightarrow$ Vehicle 1: $M_{2}:V_{1}\triangleleft\{ID_{1},T_{S},R,\mathsf{Hash}(ID_{2}\parallel Q)\}_{K_{pb1}}$ $V_{2}$ responds with its Diffie–Hellman value $R$ and a hash binding $ID_{2}$ and $Q$, encrypted with $V_{1}$’s public key.
  • Session key derivation Both vehicles, $v_{1}$, $V_{2}$ compute: $K_{s}=f(Q,R)$ using Diffie–Hellman. Subsequent messages are encrypted under $K_{s}$.
We now sketch how SG1–SG4 are derived using standard BAN rules.
  • From $M_{1}$ at $V_{2}$.
    • From $M_{1}$, $V_{2}$ receives: $V_{2}\triangleleft\{ID_{2},K_{pb1},TS,Q,Hash(ID_{2}||K_{pvt_{2}}))\}_{K_{pb2}}$
    • Using IA6 (authenticity of $K_{pb2}$) and the message meaning rule, $V_{2}$ infers: $V_{2}\mid\equiv V_{1}\mid\sim(ID_{2},K_{pb1},T_{S},Q,Hash(ID_{2}||K_{pvt_{2}}))).$
    • Using IA3 (#(Q)\#(Q)) and the nonce verification rule, $V_{2}$ derives: $V_{2}\mid\equiv V_{1}\Rightarrow Q$
    • From the belief rule, $V_{2}$ thus believes that $V_{1}$ is alive and participating in the current protocol instance. Combined with the integrity of the blockchain bindings (IA2) and consistency of $H_{2}$, $V_{2}$ concludes that the peer is the legitimate $V_{1}$, leading toward SG2:
      $V_{2}\mid\equiv V_{1}.$
  • From $M_{2}$ at $V_{1}$.
    • From $M_{2}$, $V_{1}$ receives: $V_{1}\triangleleft\{ID_{1},T_{S},R,\mathsf{Hash}(ID_{2}\parallel Q)\}_{K_{pb1}}$
    • Using IA5 (authenticity of $K_{pb1}$) and the message meaning rule, $V_{1}$ infers: $V_{1}\mid\equiv V_{2}\mid\sim(ID_{1},T_{S},R,\mathsf{Hash}(ID_{2}\parallel Q))$
    • Using IA4 (#(R)\#(R)) and the nonce verification rule, $V_{1}$ derives: $V_{1}\mid\equiv V_{2}\Rightarrow R$
    • Since the received hash $\mathsf{Hash}(ID_{2}\parallel Q)$ matches $V_{1}$’s locally computed value, $V_{1}$ is assured that $V_{2}$ correctly received and used $Q$, binding the current run of the protocol to the same peer. Together with the blockchain-confirmed credentials (IA1), this implies SG1:
      $V_{1}\mid\equiv V_{2}.$
  • Session key beliefs.
    • Both parties compute the same session key $K_{s}=f(Q,R)$. From their beliefs in the freshness of $Q$ and $R$ (IA3–IA4) and their belief that these values were correctly used by the peer, each party concludes that $K_{s}$ is fresh:
      $V_{1}\mid\equiv\#(K_{s}),$
      $V_{2}\mid\equiv\#(K_{s}).$
    • Using IA7 and IA8 (jurisdiction over shared-key statements) and the jurisdiction rule, we derive:
      $V_{1}\mid\equiv V_{1}\leftrightarrow_{K_{s}}V_{2}$ $\quad\textbf{(SG3)},$
      $V_{2}\mid\equiv V_{1}\leftrightarrow_{K_{s}}V_{2}$ $\quad\textbf{(SG4)}.$
Thus, the BAN logic derivation shows that both vehicles mutually authenticate each other and agree on a fresh session key $K_{s}$ that is known only to them, under the stated assumptions.
The BAN logic analysis confirms that the D-IM identification and authorization protocol achieves its core Zero Trust-aligned security goals: mutual authentication, fresh key establishment, and resistance to impersonation and replay, assuming the blockchain authority correctly certifies public keys and that the underlying cryptographic primitives are secure. In other words, if an attacker cannot subvert the blockchain’s credential records or break the underlying cryptography, then it cannot successfully impersonate a vehicle or derive the established session key within the modeled threat scope.
5.2 Qualitative Security Analysis
This subsection provides a qualitative security analysis of the proposed Zero Trust-based D-IM system, focusing on how the design mitigates the attacks defined in the attacker model subsection 3.3. The goal is to clearly illustrate the security rationale behind the framework and its alignment with Zero Trust principles. In line with the “never trust, always verify” philosophy, the system enforces continuous authentication, decentralized verification, and tamper-evident identity binding for all V2X entities. As a result, several attacks are prevented by design, while others are detected, contained, or their impact is significantly reduced. We also distinguish between guarantees provided by the current implementation and additional protection achievable through planned extensions such as anomaly-based detection and voting-based consensus.
  1. 1.
    Identity and Authentication Attacks
    • Masquerading / Impersonation Attacks: Impersonation attacks rely on an attacker assuming the identity of a legitimate vehicle or RSU to inject malicious messages. In the proposed D-IM system, each entity’s identity is cryptographically bound to a public key and registered on a permissioned blockchain. Mutual authentication requires cross-verification of blockchain-stored credentials before any communication is established. As a result, an attacker lacking the corresponding private key cannot successfully assume another entity’s identity. This attack is therefore prevented by design in the current implementation. Revocation records stored immutably on the blockchain further ensure that compromised identities cannot regain trust after detection.
    • Sybil Attacks: Sybil attacks attempt to introduce multiple fake identities to influence cooperative decisions. The D-IM framework mitigates this threat through decentralized identity registration and verification, where each identity must be validated by blockchain verification peers (RSUs) before being accepted. The cost and traceability of identity creation significantly raise the barrier for generating large numbers of pseudonymous identities. This provides strong prevention against large-scale Sybil attacks. Future extensions, such as behavior-based voting and reputation management, can further limit residual Sybil influence by correlating identity behavior across time.
  2. 2.
    Data Integrity and Freshness Attacks
    • Position/Velocity Data Falsification: In this attack, a legitimate but malicious participant broadcasts incorrect kinematic data. While identity verification alone cannot guarantee data correctness, the D-IM system ensures that falsified data is always attributable to a verified sender. This provides accountability and detection, rather than full prevention. Hash-based integrity checks and authenticated session keys prevent in-transit message tampering, while the Zero Trust enforcement ensures that only authenticated entities can publish messages. Planned integration with anomaly-based behavior analysis can further identify implausible trajectories or inconsistent motion patterns and limit their influence. Also, since each vehicle shares its location, during both authentication and periodic status updates, with the blockchain network residing on RSUs, RSUs can verify the reported location accuracy using the received signal strength (RSS) of corresponding received messages.
    • Replay Attacks: Replay attacks involve re-transmitting previously valid messages to mislead receivers. The proposed protocol incorporates timestamps and freshness checks during identity verification and message exchange. Messages with stale timestamps or reused cryptographic challenges are rejected during the authentication handshake. Consequently, replay attacks are prevented by design in the current implementation.
    • Sensor Spoofing (GPS, LiDAR, Radar): Sensor spoofing targets the perception layer and cannot be fully prevented by identity management alone. However, the D-IM framework limits its impact by ensuring that external data used in cooperative perception originates from authenticated and authorized entities. This enables detection and containment, as inconsistent data from authenticated senders can be flagged and correlated across multiple sources. Full mitigation is strengthened through future extensions involving multi-source validation and anomaly-based intrusion detection.
    • Man-in-the-Middle, MiTM, Attacks: MiTM attacks attempt to intercept or modify communication between legitimate entities. In the D-IM protocol, all session establishment occurs after mutual blockchain-backed identity verification and results in an encrypted session key derived via Diffie-Hellman exchange. Without access to the private keys, an intermediary cannot decrypt or alter messages without detection. As a result, active MiTM attacks are prevented, while passive eavesdropping is rendered ineffective by encryption.
  3. 3.
    Collusion and Coordinated Attacks
    • Deception and Coordinated Misinformation: Colluding vehicles may attempt to jointly broadcast false information (e.g., fake hazards). The D-IM system limits the impact of such attacks by ensuring that all data is attributable and traceable to verified identities. While collusion cannot be entirely prevented at the identity layer, the framework enables detection, attribution, and post-event accountability. Planned consensus-based validation and voting among RSUs can further reduce the influence of coordinated false reports.
    • Adversarial Machine Learning Attacks: These attacks target downstream ML models rather than the communication substrate itself. While D-IM does not directly prevent adversarial manipulation of perception models, it restricts ML inputs to authenticated data sources and preserves data integrity during transmission since only authenticated vehicles can participate in communication and data sharing. This reduces the attack surface and enables more effective detection when combined with future ML-based intrusion detection systems.
  4. 4.
    Availability and Network-Level Attacks
    • Distributed Denial-of-Service (DDoS) Attacks: DDoS attacks aim to disrupt availability by overwhelming vehicles, RSUs, or the blockchain network. The permissioned nature of the blockchain and identity-gated access control restrict participation to authenticated entities, thereby reducing the attack surface. While high-volume DDoS attacks cannot be fully prevented, the framework provides attack containment, as malicious senders can be identified, rate-limited, and revoked. Future smart-contract-based penalties can further automate this response.
  5. 5.
    Accountability and Trust Attacks
    • Repudiation Attacks: Repudiation occurs when a malicious entity denies having sent a message or performed an action. The immutable blockchain ledger records identity registration, key usage, and verification events, providing non-repudiable evidence. As a result, repudiation attacks are prevented, as entities cannot plausibly deny authenticated actions.
    • Single Point of Failure Exploitation: Traditional centralized identity systems suffer from single points of failure. By decentralizing identity verification and credential storage across multiple RSUs and blockchain peers, the D-IM framework eliminates reliance on any single trusted authority. This attack class is therefore prevented by design, with redundancy and decentralization preserving system availability even under partial compromise.
@@ 5.2 Qualitative Security Analysis
Overall, the proposed D-IM system leverages blockchain and Zero Trust principles to enable decentralized, tamper-resistant, and continuously verifiable identity management in V2X environments. The framework provides strong preventive guarantees against identity-based, replay, impersonation, man-in-the-middle, repudiation, and single-point-of-failure attacks, while enabling effective detection, attribution, and containment of data falsification, spoofing, collusion, and availability-based attacks. By ensuring rigorous authentication, efficient key establishment, and scalable operation, the proposed D-IM system establishes a robust security foundation for autonomous vehicle communication networks and successfully addresses the defined threat model. Future work will further strengthen this foundation through the integration of smart contracts and anomaly-based machine learning techniques to automate misbehavior penalties and enhance system adaptability.
6 EXPERIMENTAL EVALUATION
This section evaluates the performance and practicality of the proposed D-IM system through simulation-based experiments in realistic V2X scenarios. The evaluation focuses on quantifying the communication overhead introduced by continuous identity verification, as well as its impact on key network performance metrics under varying traffic densities. The results demonstrate that the proposed framework provides strong security and attribution guarantees while maintaining scalability and efficiency suitable for real-world autonomous vehicle deployments.
6.1 Simulation Setup
@@
$p_{k}$ (resource retention) 0.8
Sensing threshold -126 dBm
@@
This subsection presents the implementation details of our proposed identification and authorization protocol, and the simulation environment. The blockchain infrastructure is built on Hyperledger Iroha v2 [Iroha], a permissioned blockchain designed for identity-centric applications, offering a lightweight and modular architecture well suited for distributed identity management in vehicular networks. Iroha employs the Yet Another Consensus (YAC) protocol, a crash-fault-tolerant consensus mechanism that prioritizes low latency and high throughput over full Byzantine resilience. , YAC achieves eventual consistency across network nodes while maintaining system resilience in the event of node failures [Iroha_docs]. This design choice enables efficient transaction confirmation and makes Iroha a practical solution for consortium-based vehicular environments where participants are semi-trusted and real-time performance is critical.
@@
The proposed identification and authorization protocol was implemented in Python. Communication with the Iroha blockchain was facilitated through the official Iroha Python library v0.0.5.5. Cryptographic operations were realized using the cryptography package v3.4.8, which provides RSA and symmetric key primitives, while Python’s secrets module generated random secret values for Diffie-Hellman key exchange. For vehicle-to-vehicle (V2V) communication, we employed the Python socket module to establish and manage encrypted communication channels between participants.
@@
To evaluate the proposed protocol, we used WiLabV2XSim v6.2 a discrete-event simulation framework developed in MATLAB for modeling vehicular networks with a focus on cooperative awareness services. WiLabV2XSim supports both DSRC and C-V2X communications and can incorporate synthetic mobility models or real traffic traces to represent vehicle movements [bazzi2019survey]. We ran our simulations on High Performance Computing (HPC) resources supported by the University of Arizona. We dedicated one node with one core for each scenario (Different simulation scenarios explained in 6.2). Standard partition on Puma cluster was used which uses CPU only. We used 6 GBytes memory per CPU. Matlab r2023b was used to run the WiLabV2XSim simulator.
The main simulation parameters are summarized in Table 5. For DSRC, channel access is modeled using carrier-sense multiple access with collision avoidance (CSMA/CA), while C-V2X employs sensing-based semi-persistent scheduling (SPS). These settings allow evaluation of the trade-offs between communication overhead introduced by the proposed protocol and the reliability, timeliness, and channel efficiency of vehicular communication in both highway and urban environments.
6.2 Evaluation Parameters and Scenarios
Figure 5: Urban Environment Layout used
Each simulation lasted one minute and was run for both highway and urban scenarios. In the highway case, vehicles traveled along a 5000 m bidirectional road with four lanes per direction, an average speed of 120 km/h, and a standard deviation of 20 km/h. In the urban case, traffic was modeled on a $2\times 2$ grid comprising four blocks, each 250 m long and wide, with four lanes in both vertical and horizontal directions. Figure 5 shows the urban scenario layout used in our simulations. For both the highway and the urban scenarios, the lane width is 4 meters.The mean speed was 60 km/h with a standard deviation of 15 km/h. Simulations were run using DSRC and LTE-V2X communication protocols sending either the minimum sized packet or the average sized packet. We ran the simulations with different vehicle densities in both the highway and the urban scenarios. Simulations for each combination of environment, communication protocol, packet size and vehicle density were repeated 10 times with different starting seed and the outputs were averaged. The framework was evaluated using four output parameters: